OpenStudy (anonymous):

Jake cut out 3 small circles and 5 big circles from a piece of cars board. The radius of each small circle was X Inches and the radius of each big circle was Y inches. The expression below shows the total number of square inches of cardboard that jake used for cutting out the circles: 3(pie)x^2 + 5(pie)y^2 What does the term 5(pie)^2 In The Expression represent

OpenStudy (paxpolaris):

\[3\pi X^2 +5\pi Y^2 \] \(5\pi Y^2\) is the area of the 5 large circles (each with radius \(Y\))